OpenStudy (anonymous):

Will Give Medal! Which of the following is not a proportion? 4/5 = 8/10 18/6 = 12/4 6/16 = 3/8 27/9 = 22/4

OpenStudy (yrelhan4):

Which if the ratios on LHS is equal to the corresponding ratio on the RHS?

OpenStudy (anonymous):

The last one.

OpenStudy (yrelhan4):

the last ratio is true? look again..

OpenStudy (anonymous):

Oh sorry it's the second one.

OpenStudy (anonymous):

I'm sorry the third one.

OpenStudy (yrelhan4):

4/5 = 8/10.. in 8/10 both 8 and 10 are factors of 2.. 2*4=8 and 2*5=10.. so 8/10 can be written as 4/5, hence true 18/6 = 12/4 18/6=3 and 12/4=3 so this one is true as well.. 6/16 = 3/8 6/16 can be written as 3/8, as 2*3=6 and 2*8=16 so this one is true too. 27/9 = 22/4 here, 27/9=3 and 22/4=11/2 so this one isnt true..